# Hierarchical Clustering in $\Lambda$CDM Cosmologies via Persistence Energy

### 03 Jan 2024 v1

@@ Abstract
In this research, we investigate the structural evolution of the cosmic web, employing advanced methodologies from Topological Data Analysis. Our approach involves leveraging Persistence Signals, an innovative method from recent literature that facilitates the embedding of persistence diagrams into vector spaces by re-conceptualizing them as signals in +$\mathbb{R}^{2}_{+}$. Utilizing this methodology, we analyze three quintessential cosmic structures: clusters, filaments, and voids. A central discovery is the correlation between Persistence Energy and redshift values, linking persistent homology with cosmic evolution and providing insights into the dynamics of cosmic structures.
Keywords Topological Data Analysis  $\cdot$ Cosmic Web  $\cdot$ Spectral Persistence  $\cdot$ DTFE
@@ 1 Introduction
The complex geometry of persistence diagrams presents significant challenges in terms of interpretation and the extraction of meaningful insights. To address these challenges, our pipeline integrates the Spectral Persistence Homology framework, as referenced in M.E. Van Huffel (2023) and detailed in Section 5.3.1, to embed these diagrams into a vector space. Following this embedding, we examine what we term the persistence energy of the persistence diagrams, focusing on their variations across a range of redshift values from $z=5$ to $z=0$.
@@ 1.1 The Cosmic Web
Snapshots of the cosmic web capture a moment in the evolution of the large-scale structure of the universe. In the context of simulations or observational surveys, a snapshot represents a specific instance in time, typically corresponding to a particular redshift value. The redshift value associated with a snapshot is a measure of the cosmic expansion; it quantifies the extent to which the universe has stretched since the emitted light from distant objects was initially produced Planck Collaboration et al. (2020). Higher redshift values correspond to earlier cosmic times, allowing researchers to study the cosmic web at different epochs. Essentially, a snapshot with a higher redshift captures the universe when it was younger, providing insights into the formation and evolution of structures such as filaments, voids, and galaxy clusters over cosmic time scales.
One central thesis regarding the cosmic web is that the components that characterize the cosmic web, namely cluster, filaments, and voids, merge together, such that the present day web is dominated by fewer, but much more massive, structures Planck Collaboration et al. (2020).
@@ 4 Background
In their work, Chazal et al. (2013) reinterprets a persistence diagram as a measure $\mu=\sum_{\mathbf{x}\in D}m(\mathbf{x})\delta_{\mathbf{x}}$, where $D\subset\Omega$ is locally finite, and each $\mathbf{x}$ has an associated multiplicity $m(\mathbf{x})\in\mathbb{N}$ (notation from M.E. Van Huffel (2023)). Building upon this, Divol and Lacombe (2019) introduces $\mathcal{M}^{p}$, the space of Radon measures with support on $\Omega$ and a finite $p$ value, also referred to as the persistence measures space. By integrating $\mathcal{M}^{p}$ with the optimal partial transport distance, as introduced by Figalli and Gigli (2010), a metric space $(\mathcal{M}^{p},\mathrm{OT}_{p})$ is defined. [^2] The specifics of $OT_{p}$ are not detailed here as they are not crucial for this work. Unlike $\mathcal{D}_{p}$, the space $\mathcal{M}^{p}$ provides a more suitable framework for statistical analyses. This is because persistence measures are linear objects, simplifying the definition and computation of statistical quantities like means, as discussed by Divol and Lacombe (2021) and Monod (2022).
@@ 5 Methodology
The methodology presented in this paper is illustrated in Figure 1. In this pipeline, the initial input is a 3D point cloud representing dark matter subhalos. This point cloud undergoes processing through the DTFE algorithm. As described by Schaap and van de Weygaert (2000), the DTFE algorithm interpolates each point within the cloud onto a fixed-size grid, assigning a density value to each grid point. Following this, a super-level-set filtration of these density values is conducted using cubical complexes, a process efficiently detailed in the works of Wagner et al. (2011) and Mischaikow et al. (2004). Once the persistence diagram of this filtration is constructed, it is then vectorized employing methods such as Persistence Signals and persistence energy M.E. Van Huffel (2023).
@@ 5.3 Embedding of Persistence Diagrams
5.3.1 Persistence Signals
Persistence Signals represent an innovative category of descriptors within TDA, selected for their computational efficiency, theoretical grounding, and simplicity, M.E. Van Huffel (2023). This approach, considers persistence diagrams, which naturally arises as discrete measures on +$\mathbb{R}^{2}_{+}$, as signals. These diagrams are first shifted into a birth-persistence coordinate system and then transformed into a finite-dimensional vector space using a functional transform. When this functional transform maps persistence diagrams to the frequency domain, the resulting framework is referred to as Spectral Persistence Homology. For an in-depth explanation of this transformation process, we refer readers to M.E. Van Huffel (2023). In this work, we focus on two types of functional transforms: the straightforward identity transform and the more complex Gabor functional transform.
@@ 5.3.2 Persistence Energy
Persistence energy is a metric derived from the transformation coefficients of PSs. This concept is inspired by the classical notion of Fourier energy in signal processing. Analogously, the persistence energy of a measure $\mu\in\mathcal{M}^{p}$ can be mathematically expressed as:
@@
where $\kappa(\mathbf{x})$ denotes the transformation coefficient, which depends on the choice of function $f$ as specified in Equation (1). Furthermore, $\Gamma_{p}$ is defined as the space of discretized persistence diagrams (DPDs), a measure subspace of $\mathcal{M}^{p}$ in which measures are confined to allocating mass exclusively at points on a pre-defined grid. For more comprehensive details on this space, readers are referred to M.E. Van Huffel (2023).
@@
In this work, owing to the substantial heterogeneity in the multiplicity of different birth-persistence pairings in the DPDs, we precede the computation of the persistence energy of the signal induced by a measure $\nu$ with the application of a logarithmic transform on the DPDs. This transformation is essential for normalizing the varying scales of the birth-persistence pairings, thereby facilitating a more uniform and accurate analysis.
@@ 6 Results
The decrease in energy indicates that, over time, there are increasingly fewer loops created by inter-galactic filaments. The rightward shift in Figure 3 indicates that the loops are formed by filaments containing an increasing density of galaxies. We observe the same phenomenon for 0- and 2-dimensional features in Figures 5 and 6, further reinforcing the notion that the cosmic web undergoes a process of merging over time, with each characteristic structure being defined on higher density levels. Similar results are found in Wilding et al. (2021) by analysing persistence curves, a well-established vectorization of persistence diagrams.
Moreover, the DPDs of all three homology groups exhibit a high density of points near zero persistence, consistent across all redshift values. This observation is particularly relevant when examining the energy changes at different scales using Gabor transform of the DPDs (refer to Figure 7, 8, 9). At lower scales, which are sensitive to finer details, the energy values show minimal variation over $z$. This stability suggests that these fine-grained features, likely representing minor fluctuations (noise) or micro-scale patterns, remain relatively unchanged throughout the observed epochs. Since these features do not change significantly over different redshifts, it reinforces the idea that they might be inconsequential fluctuations rather than meaningful structures.
In contrast, at higher scales, there is a marked and consistent decrease in energy, indicating a substantial alteration in the larger-scale features. This drastic reduction in energy at higher scales can be attributed to the diminishing prominence of significant, high-persistence features in the cosmic web. Such features, which might represent more substantial structural formations or dense regions, appear to reduce in number or prominence as the universe evolves towards its current state, confirming the thesis that galaxies tend to merge together in an increasingly dense cosmic web Cautun et al. (2014); Planck Collaboration et al. (2020). These observations, when considered together, further reinforce the $\Lambda$CDM model of the universe.
Figure 4: log-Persistence energy for the first three homology groups, using the identity function as a transform

### 08 Jan 2024 v2

@@ Abstract
In this research, we investigate the structural evolution of the cosmic web, employing advanced methodologies from Topological Data Analysis. Our approach involves leveraging LITE, an innovative method from recent literature that embeds persistence diagrams into elements of vector spaces. Utilizing this methodology, we analyze three quintessential cosmic structures: clusters, filaments, and voids. A central discovery is the correlation between Persistence Energy and redshift values, linking persistent homology with cosmic evolution and providing insights into the dynamics of cosmic structures.
Keywords Topological Data Analysis  $\cdot$ Persistent Homology  $\cdot$ Cosmic Web  $\cdot$ $\Lambda$CDM  $\cdot$ DTFE
@@ 1 Introduction
The complex geometry of persistence diagrams presents significant interpretive challenges for extracting meaningful insights. To address this, our methodology integrates a framework for embedding these diagrams into a vector space. This approach is based on the recent work of Van Huffel and Palo (2024) and further explained in Section 5.3.1. Subsequently, we analyze what we term the persistence energy of these diagrams across varying redshift values, ranging from $z=5$ to $z=0$.
@@ 1.1 The Cosmic Web
Snapshots of the cosmic web capture a moment in the evolution of the large-scale structure of the universe. In the context of simulations or observational surveys, a snapshot represents a specific instance in time, typically corresponding to a particular redshift value. The redshift value associated with a snapshot is a measure of the cosmic expansion; it quantifies the extent to which the universe has stretched since the emitted light from distant objects was initially produced Planck Collaboration (2020). Higher redshift values correspond to earlier cosmic times, allowing researchers to study the cosmic web at different epochs. Essentially, a snapshot with a higher redshift captures the universe when it was younger, providing insights into the formation and evolution of structures such as filaments, voids, and galaxy clusters over cosmic time scales.
One central thesis regarding the cosmic web is that the components that characterize the cosmic web, namely cluster, filaments, and voids, merge together, such that the present day web is dominated by fewer, but much more massive, structures Planck Collaboration (2020).
@@ 4 Background
In their work, Chazal et al. (2013) reinterprets a persistence diagram as a measure $\mu=\sum_{\mathbf{x}\in D}m(\mathbf{x})\delta_{\mathbf{x}}$, where $D\subset\Omega$ is locally finite, and each $\mathbf{x}$ has an associated multiplicity $m(\mathbf{x})\in\mathbb{N}$ (notation from Van Huffel and Palo (2024)). Building upon this, Divol and Lacombe (2019) introduces $\mathcal{M}^{p}$, the space of Radon measures with support on $\Omega$ and a finite $p$ value, also referred to as the persistence measures space. By integrating $\mathcal{M}^{p}$ with the optimal partial transport distance, as introduced by Figalli and Gigli (2010), a metric space $(\mathcal{M}^{p},\mathrm{OT}_{p})$ is defined. [^2] The specifics of $OT_{p}$ are not detailed here as they are not crucial for this work. Unlike $\mathcal{D}_{p}$, the space $\mathcal{M}^{p}$ provides a more suitable framework for statistical analyses. This is because persistence measures are linear objects, simplifying the definition and computation of statistical quantities like means, as discussed by Divol and Lacombe (2021) and Monod (2022).
@@ 5 Methodology
The methodology presented in this paper is illustrated in Figure 1. In this pipeline, the initial input is a 3D point cloud representing dark matter subhalos. This point cloud undergoes processing through the DTFE algorithm. As described by Schaap and van de Weygaert (2000), the DTFE algorithm interpolates each point within the cloud onto a fixed-size grid, assigning a density value to each grid point. Following this, a super-level-set filtration of these density values is conducted using cubical complexes, a process efficiently detailed in the works of Wagner et al. (2011) and Mischaikow et al. (2004). Once the persistence diagram of this filtration is constructed, it is then vectorized employing methods such as LITE described in Van Huffel and Palo (2024), and persistence energy (see Section 5.3.2).
@@ 5.3 Embedding of Persistence Diagrams
5.3.1 Lattice Integrated Topological Embedding - (LITE)
LITE represent an innovative category of descriptors within TDA, selected for their computational efficiency, theoretical grounding, and simplicity, Van Huffel and Palo (2024). This approach first shift persitence diagrams into a birth-persistence coordinate system and then transform them into a elements of a finite-dimensional vector space using a functional transform. For an in-depth explanation of this transformation process, we refer readers to Van Huffel and Palo (2024). In this work, we focus on two types of functional transforms: the straightforward identity transform and the more complex Gabor functional transform.
@@ 5.3.2 Persistence Energy
Persistence energy is a metric derived from the transformation coefficients of LITE. This concept is inspired by the classical notion of Fourier energy in signal processing. Analogously, the persistence energy of a measure $\mu\in\mathcal{M}^{p}$ can be mathematically expressed as:
@@
where $\kappa(\mathbf{x})$ denotes the transformation coefficient, which depends on the choice of function $f$ as specified in Equation (1). Furthermore, $\Gamma_{p}$ is defined as the space of discretized persistence diagrams (DPDs), a measure subspace of $\mathcal{M}^{p}$ in which measures are confined to allocating mass exclusively at points on a pre-defined grid. For more comprehensive details on this space, readers are referred to Van Huffel and Palo (2024).
@@
In this work, owing to the substantial heterogeneity in the multiplicity of different birth-persistence pairings in the DPDs, we precede the computation of the persistence energy of the measure $\nu$ with the application of a logarithmic transform on the DPDs. This transformation is essential for normalizing the varying scales of the birth-persistence pairings, thereby facilitating a more uniform and accurate analysis.
@@ 6 Results
The decrease in energy indicates that, over time, there are increasingly fewer loops created by inter-galactic filaments. The rightward shift in Figure 3 indicates that the loops are formed by filaments containing an increasing density of galaxies. We observe the same phenomenon for 0- and 2-dimensional features in Figures 6 and 7, further reinforcing the notion that the cosmic web undergoes a process of merging over time, with each characteristic structure being defined on higher density levels. Similar results are found in Wilding et al. (2021) by analysing persistence curves, a well-established vectorization of persistence diagrams.
Moreover, the DPDs of all three homology groups exhibit a high density of points near zero persistence, consistent across all redshift values. This observation is particularly relevant when examining the energy changes at different scales using Gabor transform of the DPDs (refer to Figure 8, 4, 9). At lower scales, which are sensitive to finer details, the energy values show minimal variation over $z$. This stability suggests that these fine-grained features, likely representing minor fluctuations (noise) or micro-scale patterns, remain relatively unchanged throughout the observed epochs. Since these features do not change significantly over different redshifts, it reinforces the idea that they might be inconsequential fluctuations rather than meaningful structures.
In contrast, at higher scales, there is a marked and consistent decrease in energy, indicating a substantial alteration in the larger-scale features. This drastic reduction in energy at higher scales can be attributed to the diminishing prominence of significant, high-persistence features in the cosmic web. Such features, which might represent more substantial structural formations or dense regions, appear to reduce in number or prominence as the universe evolves towards its current state, confirming the thesis that galaxies tend to merge together in an increasingly dense cosmic web Cautun et al. (2014); Planck Collaboration (2020). These observations, when considered together, further reinforce the $\Lambda$CDM model of the universe.
@@
Figure 5: log-Persistence energy for the first three homology groups, using the identity function as a transform.